2x+1=-17 the answer is -9 but i want to know which is the differences between this other equation 8x+7=-3 and the answer is 2 can you explain this to me please
Answer by rfer(16322)

2x+1=-17
2x=-17-1
2x=-18
x=-18/2
x=-9
-----------
8x+7=-3
8x=-3-7
8x=-10
x=-10/8
x=-1.25
-----------
I don't know where you got 2 from?
Both equations have the same operations.